OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <6, -1>, v = <7, -4>

OpenStudy (dumbcow):

for any 2 vectors, the following relationship holds \[\cos \theta = \frac{u*v}{|u|*|v|}\]

OpenStudy (anonymous):

so how would i proceed with this problem?

OpenStudy (dumbcow):

find magnitudes of u and v , find dot product of u and v then solve for theta

OpenStudy (anonymous):

could you please show me how to do that step by step? :)

OpenStudy (dumbcow):

dot product of 2 vectors \[u = <a,b> , v=<c,d>\] \[u*v = (a*c)+(b*d)\] magnitude of a given vector \[u = <a,b>\] \[|u| = \sqrt{a^{2} +b^{2}}\]

OpenStudy (anonymous):

i got 8.06 and 6.08

OpenStudy (dumbcow):

for what ?

OpenStudy (anonymous):

the magintude of the vectors?

OpenStudy (dumbcow):

ahh yes correct....i left them as sqrt(37) and sqrt(65)

OpenStudy (dumbcow):

be careful with rounding since that can affect the final answer

OpenStudy (anonymous):

what is the angle between them? :)

OpenStudy (dumbcow):

ok what is dot product (u*v) ?

OpenStudy (anonymous):

42

OpenStudy (dumbcow):

not quite...should get 46 --> 7*6 + (-1)*(-4) = 42+4 = 46

OpenStudy (dumbcow):

\[\cos \theta = \frac{46}{\sqrt{37} \sqrt{65}}\] take inverse cos (use calculator) \[\theta = \cos^{-1} \frac{46}{\sqrt{37} \sqrt{65}} \approx 20.3^{o}\]

OpenStudy (anonymous):

thank you :)